A)A solenoid is made of 3000 turns of copper wire. The length of the coil is 35.0cm. It is connected to a power supply which provides a current of 5.00A. What is the magnitude of the B-field created inside the solenoid.
B) The solenoid coil is twisted to make a toroid of radius 10.0cm. Determine the magnetic field magnitude inside the coils of the toroid.
Answer #1

A) B-field = uoNI / L

B-field = 4*pi*10-7 * 3000*5 / 0.35

B-field = 0.05385 T

B)  B-field = uoNI / 2*pi*r

B-field = 4*pi*10-7 * 3000*5 / 2*pi*0.10

B-field = 0.03 T